OpenStudy (sleepyjess):

If 48.5 grams of aluminum metal (Al) react with 62.8 grams of sulfur (S) in a synthesis reaction, how many grams of the excess reactant will be left over when the reaction is complete?

OpenStudy (sleepyjess):

I know how to balance an equation and I know how to compute everything, but what I don't understand is how to balance this equation since all they've given me is how many grams of each.

OpenStudy (sleepyjess):

@perl

OpenStudy (sleepyjess):

@SolomonZelman @TuringTest

OpenStudy (sleepyjess):

@Jhannybean

OpenStudy (perl):

can you write out the balanced equation first

OpenStudy (sleepyjess):

That's what I don't understand how to do because that is all the information they have given.

OpenStudy (sleepyjess):

I found this, but I still don't understand the explanation: A synthesis reaction is the one in which a particular compound is synthesized by the chemical reaction. The question you've asked is also an example of synthesis reaction between Al and S to yield the product Al2S3. To solve such questions, first of all, write down the balanced chemical equation of the reaction. 2Al + 3S ---> Al2S3 Given amount of Al = 48.5g => no. of moles of Al present = given amount/atomic mass = 48.5/27 = 1.796 mol Given amount of S = 62.8g => no. of moles of S present = given amount/atomic mass = 62.8/32 = 1.963 mol But according to the balanced chemical equation, Al and S react in the ratio of 2:3. => 1 mol of Al will react with 3/2 mol of S. => 1.796 mol of Al will react with (3/2)* 1.796 mol of S, i.e., 2.694 mol of S. But actually, the amount of S present is less than this value. => S is the limiting reagent, or AL is the excess reagent. Now, 1 mol of S will react with 2/3 mol of Al. => 1.963 mol of S will react with (2/3)*1.963 mol or 1.309 mol of Al. No. of moles of Al left = 1.796 - 1.309 = 0.487 mol Amount of Al left = moles * Atomic mass = 0.487 * 27 = 13.15g

OpenStudy (sleepyjess):

Where did they get 2 Al + 3S --> Al2S3?

OpenStudy (jhannybean):

Aluminum and sulfur react.

OpenStudy (jhannybean):

Combination reaction, so they combine to form aluminum sulfide

OpenStudy (jhannybean):

\[\sf 2Al (s) + 3S (g) \longrightarrow Al_2S_3(s) \]Given: \(\sf \text{48.5 g Al}\) and \(\sf \text{62.8 g S}\) \[\sf 48.5 ~\text{g Al} \cdot \frac{1~\text{mol Al}}{26.98~\text{g Al}}= ~?\]

OpenStudy (sleepyjess):

How do you know that? This is the first time that they haven't given us an equation to balance. Is aluminum sulfide the only thing that it can create?

OpenStudy (perl):

I think by 'synthesis reaction' they mean they combine to make an ionic or molecular substance

OpenStudy (jhannybean):

Because you need to know elements combine with eachother, and how they react according to the types of metals they are, and what they produce.

OpenStudy (sleepyjess):

I know how to do the rest of it, I'm just fighting myself on how you knew 2Al(s)+3S(g)⟶Al2S3(s)

OpenStudy (jhannybean):

You have learned -single replacement -double replacement - combination -decomposition -combustion havent you?

OpenStudy (sleepyjess):

Yes

OpenStudy (perl):

aluminum is a metal , and sulfur is a non metal, so they combine to make an ionic compound

OpenStudy (jhannybean):

Exactly :)

OpenStudy (sleepyjess):

Ok, thank you perl and jhanny

OpenStudy (jhannybean):

are you good with the rest of it?

OpenStudy (sleepyjess):

Yep

OpenStudy (jhannybean):

Ok then :)

OpenStudy (anonymous):

No.. This is not a practice exam. This is a module quiz from FLVS and it's cheating if you use this answer. Plus it has a obvious mistake in it. Mr. Carlyle FLVS Chemistry Instructor This question is a violation of the OpenStudy Guidelines. The question is from an online school plus a violation by solicitation for members to unknowingly assist them cheating on an exam. 30653065chem30653065

OpenStudy (anonymous):

lol

sammixboo (sammixboo):

@carlyle She told me this "It actually was on the practise quiz, but I didn't know that when I asked it. I searched one of the practice problems in the lesson and that is what came up. I didn't understand older posts, so I asked it"

OpenStudy (anonymous):

You're cool for asking because that's how we learn from one another. The goal is learning how to work the problem so you're able to work any other similar one. You're going to have your own symbols, calculations, work steps along with trying to make the answer a little better like following the sig fig rules.

OpenStudy (sleepyjess):

Just wondering, what is this? 30653065chem30653065

OpenStudy (anonymous):

Oh that's a tag, so when I search for it the unique set of characters I'm able to find my responses in case I should remove or modify them.